There are three bus lines between X and Y; and five bus lines between Y and Z. Find the number of ways a person can travel Round trip by bus from X to Z by way of Y, if the person does not want to use a bus line more than once.

Answer by ikleyn(52776)

.
(3*5)*(4*2) = 15*8 = 120.